OpenStudy (anonymous):

w(a+6)+2w working out and answer please

OpenStudy (aripotta):

do you know what you have to do first?

OpenStudy (anonymous):

No

OpenStudy (amistre64):

it might be good to know what defines a "solution"

OpenStudy (aripotta):

you have to distribute the w into the parentheses

OpenStudy (aripotta):

and i think it just needs to be simplified

OpenStudy (amistre64):

Ari, we could also factor out a "w" from the expression ;)

OpenStudy (amistre64):

w(a+6)+2w w(a+6+2)

OpenStudy (aripotta):

but that wouldn't be simplified, would it?

OpenStudy (anonymous):

There are four questions on this sheet it says expand the following and simplify if possible. Grade 7 algebra of my daughters I am trying to work it out so I can help her.

OpenStudy (amistre64):

it could be a step towards simplification but if the sheet says to expand first, then yes; distribution would be more proper

OpenStudy (amistre64):

given a general setup, distribution (expanding) takes the form \[a(b+c)\to\ ab+ac\]

OpenStudy (anonymous):

distribute first so its goin to end up as wa+6w+2w then you combine like terms so combine w's first 9w+a now the variables are =1 9(1)+(1) =9+1=10

OpenStudy (amistre64):

you can view it from the perspective of just mulitplying numbers in the vertical setup b + c a ------ ab + ac this is useful for more multi termed constructions

OpenStudy (amistre64):

jay, your workings are a little off

OpenStudy (aripotta):